# Cross-lingual Offensive Language Detection: A Systematic Review of Datasets, Transfer Approaches and Challenges

### v1

@@ 4. Multilingual Hate Speech Datasets
EXIST@IberLEF (Rodríguez-Sánchez et al., 2021) |
Sexism Identification in Social Networks |
2021 |
sexism |
en, es |
31 |
UrduThreat@FIRE (Amjad et al., 2022) |
@@
Abusive and Threatening Language Detection in Urdu |
2021 |
abuse, threat |
ur |
19 |
@@
Moj Multilingual Abusive Comment Identification across Indic Languages |
2021 |
abuse |
10+ Indic |
54 | |
HASOC@FIRE (Satapara et al., 2023) |
@@
Offensive Content Identification in English and Indo-Aryan Languages |
2022 |
hate speech, offense |
en, hi, de, mr |
12 |
@@
EXIST@IberLEF (Rodríguez-Sánchez et al., 2022) |
Sexism Identification in Social Networks |
2022 |
sexism |
en, es |
19 |
HaSpeeDe@Evalita (Lai et al., 2023) |
Political and Religious Hate Speech Detection |
2023 |
hate speech |
it |
6 |
EXIST@IberLEF (Plaza et al., 2023) |
Sexism Identification in Social Networks |
2023 |
sexism |
en, es |
28 |
@@
Hate Speech and Offensive Content Identifica- |
2023 |
hate speech, |
en, hi, de, |
tion in English and Indo-Aryan Languages |
offense |
Indo-Aryan |
@@ 5. Cross-lingual Resources
5.1. Multilingual linguistic Resources
There are some basic but essential linguistic resources that facilitate the task of cross-lingual hate speech detection across two or more languages. Among these, multilingual lexicons and parallel corpora stand out as foundational resources, frequently utilised by researchers to bridge linguistic gaps and enhance model performance in CLTL.
@@ 5.1. Multilingual linguistic Resources
Due to the lack of sufficient labelled datasets for hate speech, multilingual lexicons can help bridge this gap by providing connections between resource-rich and resource-poor languages (Pamungkas et al., 2021a; Jiang and Zubiaga, 2021), while parallel corpora provide translations of labelled instances from resource-rich to resource-poor languages, to facilitate CLTL (Rivaling, 2021; Stamou et al., 2022; Bigoulaeva et al., 2023). Cross-lingual word embeddings are also created by utilising these multilingual resources. In addition, these are also valuable linguistic resources in other NLP tasks, such as machine translation and multilingual model training.
Multilingual linguistic resources are most often created for specific domains, such as hate speech and abusive language. While hate speech patterns can be universal, they can also be specific to certain cultures or languages. These resources provide insights into how hate speech manifests differently across languages and cultures, furthering the understanding of cultural nuances (Nozza, 2021).
5.2. Machine Translation and Transliteration Tools
Translation refers to converting content from one language into another, whereas transliteration is the process of converting text from one script into another, in both cases preserving the meaning of the original text (Hande et al., 2021b). Machine translation tools can translate datasets from one language to another, such as Google Translate,666https://translate.google.com Microsoft Translator,777https://translator.microsoft.com/ DeepL,888https://www.deepl.com/en/translator) and machine translation models (e.g., mBART (Liu et al., 2020) and mT5 (Xue et al., 2021)). Besides, machine transliteration tools are essential for code-mixed datasets, such as Google Transliteration,999https://www.google.co.in/inputtools/try/ Microsoft Transliteration,101010https://www.microsoft.com/en-us/translator/business/translator-api/ AI4Bharat transliteration application,111111https://github.com/AI4Bharat/IndianNLP-Transliteration and transliteration python packages (e.g., indic-transliteration121212https://github.com/indic-transliteration/indic_transliteration_py).
@@ 5.2. Machine Translation and Transliteration Tools
However, in offensive language detection, some translation tools may lead to translation errors or cultural nuances that are inconsistent with the original meaning (Pamungkas et al., 2021a). For instance, derogatory terms or slurs without direct equivalents in other languages, or their severity might differ across cultures. Additionally, idiomatic expressions that convey hate or abuse in one language might lose their offensive meaning when translated literally. Therefore, translation quality is crucial to maintain the effectiveness and robustness of cross-lingual detection models.
5.3. Multilingual Representations
@@
Multilingual representations, as language independent representations, are often used in cross-lingual transfer learning. With these representations we can directly address the difference between source and target languages by projecting the text into a shared feature space. Then it is easier to project the behaviour of the model. In general, we can distinguish between using word-level representations and sentence-level representations.
@@
5.4. Monolingual and Multilingual Pre-trained Language Models
Pre-trained language models (PLMs) are a SOTA technique in NLP, as a language model trained on large amounts of data. PLMs can be initialised and further trained or fine-tuned with target data for cross-lingual transfer learning. The best-known examples of PLMs are Embeddings from Language Models (ELMo) (Peters et al., 2018) and Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019).
Multilingual PLMs are an extension of monolingual PLMs, trained on multiple languages at the same time. They learn to understand the connections between languages by developing a shared linguistic representation. There is no need to provide multilingual PLMs with any additional information about interlanguage relations, and cross-lingual transfer is possible between any languages (Pikuliak et al., 2021). Prominent examples include Multilingual BERT (mBERT) (Devlin et al., 2019) and XLM-RoBERTa (XLM-R) (Conneau et al., 2020), which have demonstrated remarkable capabilities in understanding and processing multiple languages simultaneously.
5.5. Language-Agnostic Resources
Language-agnostic resources (e.g., emojis, capitals, and punctuations) can be seen as common traits among different languages, sharing similar knowledge to enhance the semantic connection across languages. For instance, emojis are generally associated with emotional expressions, which in turn are associated with various forms of online hate (Corazza et al., 2020).
@@ 6. Cross-lingual Transfer Approaches
In this section, we systematically describe and compare different cross-lingual techniques to offensive language online. Referring to the way of identifying different transfer learning techniques in Pan and Yang (2010) and Pikuliak et al. (2021), we analyse cross-lingual approaches in all reviewed papers and ultimately decide to categorise them according to “what to transfer” – the level at which knowledge transfer or sharing occurs. The overall hierarchy of various cross-lingual transfer approaches is depicted in Figure 4. Cross-lingual transfer paradigms consist of three main categories:
@@
We look at transfer levels in Table 4 and models in Table 5. Additionally, a comprehensive Table is provided online131313https://github.com/aggiejiang/crosslingual-offensive-language-survey to summarise all techniques utilised in these studies.
@@
(Mathur et al., 2018b; Ibrohim and Budi, 2019a; Chiril et al., 2019; Pamungkas and Patti, 2019; Pant and Dadu, 2020; Arango et al., 2022; Elouali et al., 2020; Pamungkas et al., 2020; Aluru et al., 2020; Jiang and Zubiaga, 2021; Song et al., 2021; Hande et al., 2021b; Vashistha and Zubiaga, 2021; Pamungkas et al., 2021a; Aluru et al., 2021; Salaam et al., 2022; Ryzhova et al., 2022; Boulouard et al., 2022; Shi et al., 2022; Deshpande et al., 2022; Zia et al., 2022; Das et al., 2022; zahra El-Alami et al., 2022; Kumar and Sachdeva, 2022; Zhou et al., 2023) |
Feature |
@@
(Sohn and Lee, 2019; Corazza et al., 2020; Bansal et al., 2020; Aluru et al., 2020; Arango et al., 2021; Hahn et al., 2021; Rodríguez et al., 2021; Vitiugin et al., 2021; Pelicon et al., 2021a; Bigoulaeva et al., 2021, 2022; De la Peña Sarracén and Rosso, 2022; Shi et al., 2022; Jiang and Zubiaga, 2023; Bigoulaeva et al., 2023; Paul et al., 2023) |
Parameter |
@@
(Mathur et al., 2018b; Chiril et al., 2019; Kapoor et al., 2019; Sohn and Lee, 2019; Pamungkas and Patti, 2019; Ousidhoum et al., 2019; Pant and Dadu, 2020; Arango et al., 2022; Elouali et al., 2020; Stappen et al., 2020; Glavaš et al., 2020; Pamungkas et al., 2020; Ranasinghe and Zampieri, 2020; Aluru et al., 2020; Markov et al., 2021; Burtenshaw and Kestemont, 2021; Firmino et al., 2021; Tiţa and Zubiaga, 2021; Jiang and Zubiaga, 2021; Rodríguez et al., 2021; Song et al., 2021; Pelicon et al., 2021b, a; Bigoulaeva et al., 2021; Ranasinghe and Zampieri, 2021a; Hande et al., 2021a, b; Ranasinghe and Zampieri, 2021b; Vashistha and Zubiaga, 2021; Ranasinghe and Zampieri, 2021c; Nozza, 2021; Gaikwad et al., 2021; Pamungkas et al., 2021a; Aluru et al., 2021; Salaam et al., 2022; Husain and Uzuner, 2022; Gupta et al., 2022; Montariol et al., 2022; Ryzhova et al., 2022; Muti and Barrón-Cedeño, 2022; Phung and Cloos, 2022; Bigoulaeva et al., 2022; Boulouard et al., 2022; R”̈ottger et al., 2022a; De la Peña Sarracén and Rosso, 2022; Shi et al., 2022; Deshpande et al., 2022; Eronen et al., 2022; Mozafari et al., 2022; Zia et al., 2022; R”̈ottger et al., 2022b; Das et al., 2022; zahra El-Alami et al., 2022; Kumar and Sachdeva, 2022; Plaza-del arco et al., 2023; Zhou et al., 2023; Bigoulaeva et al., 2023; Awal et al., 2023) |
Type of Model | Model | Reference |
Machine Learning |
NB |
(Ibrohim and Budi, 2019a) |
Based |
LR |
(Aluru et al., 2020; Vashistha and Zubiaga, 2021; Aluru et al., 2021; Arango Monnar et al., 2022; Deshpande et al., 2022; Paul et al., 2023) |
SVM |
(Ibrohim and Budi, 2019a; Pamungkas and Patti, 2019; Pamungkas et al., 2020; Jiang and Zubiaga, 2021; Rodríguez et al., 2021; Bigoulaeva et al., 2021; Paul et al., 2023) |
DT | ||
RF |
@@
(Ibrohim and Budi, 2019a; Burtenshaw and Kestemont, 2021; Rodríguez et al., 2021) |
other |
(Hahn et al., 2021) |
Deep Learning |
MLP |
(Jiang and Zubiaga, 2021; Pelicon et al., 2021a; Kumar and Sachdeva, 2022; Paul et al., 2023) |
Based |
CNN |
@@
(Mathur et al., 2018b; Elouali et al., 2020; Markov et al., 2021; Jiang and Zubiaga, 2021; Bigoulaeva et al., 2021, 2022; Jiang and Zubiaga, 2023; Bigoulaeva et al., 2023; Paul et al., 2023) |
LSTM |
@@
(Chiril et al., 2019; Kapoor et al., 2019; Pamungkas and Patti, 2019; Ousidhoum et al., 2019; Arango et al., 2022; Stappen et al., 2020; Pamungkas et al., 2020; Markov et al., 2021; Burtenshaw and Kestemont, 2021; Jiang and Zubiaga, 2021; Vitiugin et al., 2021; Bigoulaeva et al., 2021; Pamungkas et al., 2021a; Phung and Cloos, 2022; Bigoulaeva et al., 2022; Shi et al., 2022; Arango Monnar et al., 2022; Kumar and Sachdeva, 2022; Bigoulaeva et al., 2023; Paul et al., 2023) |
GNN |
(De la Peña Sarracén and Rosso, 2022) |
other |
@@
(Ousidhoum et al., 2019; Bansal et al., 2020; Jiang and Zubiaga, 2021; Vashistha and Zubiaga, 2021; Arango Monnar et al., 2022; Deshpande et al., 2022; Kumar and Sachdeva, 2022) |
Transformer Based |
BERT |
@@
(Sohn and Lee, 2019; Pamungkas et al., 2020; Aluru et al., 2020; Markov et al., 2021; Arango et al., 2021; Hande et al., 2021a; Vashistha and Zubiaga, 2021; Muti and Barrón-Cedeño, 2022; Boulouard et al., 2022; Arango Monnar et al., 2022; Zia et al., 2022; zahra El-Alami et al., 2022; Plaza-del arco et al., 2023; Zhou et al., 2023; Jiang and Zubiaga, 2023; Paul et al., 2023) |
RoBERTa |
(Pant and Dadu, 2020; Hande et al., 2021a; Zia et al., 2022; Plaza-del arco et al., 2023; Zhou et al., 2023) |
DistilBERT |
(Hande et al., 2021a) |
AraBERT |
(Husain and Uzuner, 2022; Boulouard et al., 2022; zahra El-Alami et al., 2022) |
ULMFiT | ||
mBERT |
@@
(Sohn and Lee, 2019; Glavaš et al., 2020; Aluru et al., 2020; Burtenshaw and Kestemont, 2021; Tiţa and Zubiaga, 2021; Jiang and Zubiaga, 2021; Rodríguez et al., 2021; Song et al., 2021; Pelicon et al., 2021b, a; Hande et al., 2021b; Ranasinghe and Zampieri, 2021c; Nozza, 2021; Pamungkas et al., 2021a; Aluru et al., 2021; Montariol et al., 2022; Muti and Barrón-Cedeño, 2022; Bigoulaeva et al., 2022; Boulouard et al., 2022; De la Peña Sarracén and Rosso, 2022; Deshpande et al., 2022; Eronen et al., 2022; Das et al., 2022; zahra El-Alami et al., 2022; Zhou et al., 2023; Bigoulaeva et al., 2023; Awal et al., 2023) |
XLM |
(Corazza et al., 2020; Stappen et al., 2020; Hande et al., 2021a; Zhou et al., 2023) |
XLM-R |
@@
(Pant and Dadu, 2020; Glavaš et al., 2020; Ranasinghe and Zampieri, 2020; Firmino et al., 2021; Tiţa and Zubiaga, 2021; Jiang and Zubiaga, 2021; Song et al., 2021; Ranasinghe and Zampieri, 2021a; Hande et al., 2021a, b; Ranasinghe and Zampieri, 2021b, c; Nozza, 2021; Gaikwad et al., 2021; Salaam et al., 2022; Gupta et al., 2022; Montariol et al., 2022; Ryzhova et al., 2022; De la Peña Sarracén and Rosso, 2022; Arango Monnar et al., 2022; Eronen et al., 2022; Mozafari et al., 2022; Zia et al., 2022; Plaza-del arco et al., 2023; Awal et al., 2023) |
XLM-T | ||
MuRIL | ||
other |
@@
(Rodríguez et al., 2021; Pelicon et al., 2021b; Hande et al., 2021a, b; Husain and Uzuner, 2022; Gupta et al., 2022; Montariol et al., 2022; Plaza-del arco et al., 2023; Jiang and Zubiaga, 2023) |
Large Language |
FLAN-T5 |
(Plaza-del arco et al., 2023) |
Model |
mT0 |
(Plaza-del arco et al., 2023) |
6.1. Instance Transfer
Instances in the cross-lingual hate speech task are comprised of two elements – texts and corresponding labels – in both source and target languages. Although the single source language data is not directly applicable in cross-lingual transfer, there are certain parts of the instance that can still be repurposed together with target language data via instance transfer. Instance transfer (also called instance projection) revolves around the transfer of specific data elements (either text or label information) between source and target languages, which is a key technique in the realm of cross-lingual offensive language detection.
@@ 6.1. Instance Transfer
To implement instance transfer, a correspondence between instances in the source and target languages must be created on the data level. Correspondence refers to the pair of instances with the same meaning of texts or identical labels (Pikuliak et al., 2021), e.g., parallel data possess the same label matching, and translated text retains the same label of its original counterpart. Based on the established correspondence, texts or labels can be transitioned from one language to the other . Subsequently, these projected texts or labels can be used for further training and fine-tuning stages.

### v2

@@ 4. Multilingual Hate Speech Datasets
| EXIST@IberLEF (Rodr´ıguez-Sánchez et al., 2021) | Sexism Identification in Social Networks | 2021 | sexism | en, es | 31 |
@@
| UrduThreat@FIRE (Amjad et al., 2022) | Abusive and Threatening Language Detection in Urdu | 2021 | abuse, threat | ur | 19 |
@@
| IIIT-D@Kaggle | Moj Multilingual Abusive Comment Identification across Indic Languages | 2021 | abuse | 10+ Indic | 54 |
@@
| HASOC@FIRE (Satapara et al., 2023) | Offensive Content Identification in English and Indo-Aryan Languages | 2022 | hate speech, offense | en, hi, de, mr | 12 |
@@
| EXIST@IberLEF (Rodr´ıguez-Sánchez et al., 2022) | Sexism Identification in Social Networks | 2022 | sexism | en, es | 19 |
| HaSpeeDe@Evalita (Lai et al., 2023) | Political and Religious Hate Speech Detection | 2023 | hate speech | it | 6 |
| EXIST@IberLEF (Plaza et al., 2023) | Sexism Identification in Social Networks | 2023 | sexism | en, es | 28 |
@@
| HASOC@FIRE | Hate Speech and Offensive Content Identifica- | 2023 | hate speech, offense | en, hi, de, | - |
| tion in English and Indo-Aryan Languages | Indo-Aryan |
Competitions and Shared Tasks. The rising concern surrounding online abuse has led to the establishment of numerous competitions and shared tasks within both national (e.g., GermEval, Evalita, IberLEF) and international (e.g., SemEval) evaluation campaigns. These open scientific competitions release benchmark datasets and invite participants to submit detection results and detailed system reports (Poletto et al., 2021). Table 3 lists 24 such competitions and shared tasks from 2018 to 2023, each addressing different facets of offensive content across multiple languages.
Prominent topics include offense and HS. Specifically, GermEval competitions in 2018 (Wiegand et al., 2018) and 2019 (Strus̈ et al., 2019) centre around identifying offensive content in German tweets. OffensEval, initially in 2019 focused only on English content (Zampieri et al., 2019), expanded to include multiple languages in the subsequent edition in 2020 (Zampieri et al., 2020). The task in OSACT4 (Mubarak et al., 2020), the 4th Workshop on Open-Source Arabic Corpora and Processing Tools, focuses on offensive language detection in Arabic, especially its dialectal forms. HaSpeeDe has consistently addressed HS with various general and particular topics (Bosco et al., [n. d.]; Sanguinetti et al., 2020; Lai et al., 2023). For instance, HaSpeeDe (Bosco et al., [n. d.]) and HaSpeeDe 2 (Sanguinetti et al., 2020) identify HS against immigrants and Muslims, while HaSpeeDe 3 (Lai et al., 2023) explores HS in strongly polarised debates, particularly political and religious topics. Similarly, the HASOC series, recurring annually from 2019 to 2023, identify HS and offensive content across various languages (Mandl et al., 2019, 2021; Modha et al., 2021; Satapara et al., 2023). TRAC workshops, held in 2018 (Kumar et al., 2018) and 2020 (Kumar et al., 2020), tackled trolling, aggression, and cyberbullying. Other competitions, like PolEval (Ptaszynski et al., 2019) and UrduThreat (Amjad et al., 2022), focus on specific languages (such as Polish and Urdu) or topics (e.g., cyberbullying and threatening languages). The Jigsaw Multilingual Toxic Comment Classification competition on Kaggle attracted over 1600 participants to build multilingual models using English-only training data. Another Kaggle competition, IIIT-D Multilingual Abusive Comment Identification, focuses on abusive comments in Indic languages. Additionally, some tasks focus on more specific topics, such as sexism and misogyny. AMI is organised by IberEval (Fersini et al., 2018b) and Evalita (Fersini et al., 2018a) in 2018, specifically addressing Spanish and Italian misogynous content respectively. EXIST is held for three consecutive years from 2021 to 2023 and aims to identify sexism in social networks (Rodr´ıguez-Sánchez et al., 2021, 2022; Plaza et al., 2023). Most of these benchmark datasets originate from social media platforms, and often conduct annotations that go beyond simple binary labels to finer-grained ones.
@@ 5. Cross-lingual Resources
This section reviews the cross-lingual resources that enable offensive language detection, including datasets, lexicons, models, and auxiliary tools that facilitate transfer across languages.
Multilingual linguistic Resources. There are some basic but essential linguistic resources that facilitate the task of cross-lingual HS detection across two or more languages. Among these, multilingual lexicons and parallel corpora stand out as foundational resources, frequently utilised by researchers to bridge linguistic gaps and enhance model performance in CLTL.
@@
Due to the lack of sufficient labelled datasets for HS, multilingual lexicons can help bridge this gap by providing connections between resource-rich and resource-poor languages (Pamungkas et al., 2021a; Jiang and Zubiaga, 2021), while parallel corpora provide translations of labelled instances from resource-rich to resource-poor languages, to facilitate CLTL (Rivaling, 2021; Stamou et al., 2022; Bigoulaeva et al., 2023). Cross-lingual word embeddings are also created by utilising these multilingual resources. These resources are most often created for specific domains, such as HS and abusive language. While HS patterns can be universal, they can also be specific to certain cultures or languages. They provide insights into how HS manifests differently across languages and cultures, furthering the understanding of cultural nuances (Nozza, 2021).
@@
Machine Translation and Transliteration Tools. Translation refers to converting content from one language into another, whereas transliteration is the process of converting text from one script into another, in both cases preserving the meaning of the original text (Hande et al., 2021b). Machine translation tools can translate datasets from one language to another, such as Google Translate,666https://translate.google.com Microsoft Translator,777https://translator.microsoft.com/ DeepL888https://www.deepl.com/en/translator) and machine translation models (e.g., mBART (Liu et al., 2020) and mT5 (Xue et al., 2021)). Besides, machine transliteration tools are essential for code-mixed datasets, such as Google Transliteration,999https://www.google.co.in/inputtools/try/ Microsoft Transliteration,101010https://www.microsoft.com/en-us/translator/business/translator-api/ AI4Bharat transliteration application,111111https://github.com/AI4Bharat/IndianNLP-Transliteration and transliteration python packages (e.g., indic-transliteration121212https://github.com/indic-transliteration/indic_transliteration_py). However, in offensive language detection, some translation tools may lead to translation errors or cultural nuances that are inconsistent with the original meaning (Pamungkas et al., 2021a). For instance, derogatory terms or slurs without direct equivalents in other languages, or their severity might differ across cultures. Additionally, idiomatic expressions that convey abuse in one language might lose their offensive meaning when translated literally. Therefore, translation quality is crucial to maintain the effectiveness and robustness of cross-lingual models.
@@
Multilingual Representations. Multilingual representations, as language independent representations, are often used in cross-lingual transfer learning. With these representations we can directly address the difference between source and target languages by projecting the text into a shared feature space. Then it is easier to project the behaviour of the model. In general, we can distinguish between using word-level representations and sentence-level representations.
@@
Monolingual and Multilingual Pre-trained Language Models. Pre-trained language models (PLMs) are a SOTA technique in NLP, as a language model trained on large amounts of data. PLMs can be initialised and further trained or fine-tuned with target data for cross-lingual transfer learning. The best-known examples of PLMs are Embeddings from Language Models (ELMo) (Peters et al., 2018) and Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019). Multilingual PLMs are an extension of monolingual PLMs, trained on multiple languages. They learn the connections between languages by developing a shared linguistic representation. There is no need to provide multilingual PLMs with any additional information about interlanguage relations, and cross-lingual transfer is possible between any languages (Pikuliak et al., 2021). Prominent models include Multilingual BERT (mBERT) (Devlin et al., 2019) and XLM-RoBERTa (XLM-R) (Conneau et al., 2020), demonstrating remarkable capabilities in understanding and processing multiple languages simultaneously.
Language-Agnostic Resources Language-agnostic resources (e.g., emojis, capitals, and punctuations) can be seen as common traits among different languages, sharing similar knowledge to enhance the semantic connection across languages. For instance, emojis are generally associated with emotional expressions, which in turn are associated with various forms of online hate (Corazza et al., 2020).
@@ 6. Cross-lingual Transfer Approaches
In this section, we systematically describe and compare different cross-lingual techniques for offensive language. Referring to the way of identifying different transfer learning techniques in Pan and Yang (2010) and Pikuliak et al. (2021), we analyse cross-lingual approaches in all reviewed papers and ultimately decide to categorise them according to “what to transfer” – the level at which knowledge transfer or sharing occurs. The overall hierarchy of various cross-lingual transfer approaches is depicted in Figure 4. Cross-lingual transfer paradigms consist of three main categories:
@@
We summarise studies based on transfer levels in Table 4 and model types in Table 5. Additionally, a comprehensive Table is provided online131313https://github.com/aggiejiang/crosslingual-offensive-language-survey to summarise all techniques utilised in these studies.
6.1. Instance Transfer
Instances in the cross-lingual HS task are comprised of two elements (texts and corresponding labels) in both source and target languages. Although the single source language data is not directly applicable in cross-lingual transfer, there are certain parts of the instance that can still be repurposed together with target language data via instance transfer. Instance transfer (or instance projection) revolves around the transfer of specific data elements (either text or label information) between source and target languages, which is a key technique in the realm of cross-lingual HS detection.
@@ 6.1. Instance Transfer
| Instance | (Mathur et al., 2018b; Ibrohim and Budi, 2019a; Chiril et al., 2019; Pamungkas and Patti, 2019; Pant and Dadu, 2020; Arango et al., 2022; Elouali et al., 2020; Pamungkas et al., 2020; Aluru et al., 2020; Jiang and Zubiaga, 2021; Song et al., 2021; Hande et al., 2021b; Vashistha and Zubiaga, 2021; Pamungkas et al., 2021a; Aluru et al., 2021; Salaam et al., 2022; Ryzhova et al., 2022; Boulouard et al., 2022; Shi et al., 2022; Deshpande et al., 2022; Zia et al., 2022; Das et al., 2022; zahra El-Alami et al., 2022; Kumar and Sachdeva, 2022; Zhou et al., 2023) |
@@
| Feature | (Sohn and Lee, 2019; Corazza et al., 2020; Bansal et al., 2020; Aluru et al., 2020; Arango et al., 2021; Hahn et al., 2021; Rodríguez et al., 2021; Vitiugin et al., 2021; Pelicon et al., 2021a; Bigoulaeva et al., 2021, 2022; De la Peña Sarracén and Rosso, 2022; Shi et al., 2022; Jiang and Zubiaga, 2023; Bigoulaeva et al., 2023; Paul et al., 2023) |
@@
| Parameter | (Mathur et al., 2018b; Chiril et al., 2019; Kapoor et al., 2019; Sohn and Lee, 2019; Pamungkas and Patti, 2019; Ousidhoum et al., 2019; Pant and Dadu, 2020; Arango et al., 2022; Elouali et al., 2020; Stappen et al., 2020; Glavaš et al., 2020; Pamungkas et al., 2020; Ranasinghe and Zampieri, 2020; Aluru et al., 2020; Markov et al., 2021; Burtenshaw and Kestemont, 2021; Firmino et al., 2021; Tiţa and Zubiaga, 2021; Jiang and Zubiaga, 2021; Rodríguez et al., 2021; Song et al., 2021; Pelicon et al., 2021b, a; Bigoulaeva et al., 2021; Ranasinghe and Zampieri, 2021a; Hande et al., 2021a, b; Ranasinghe and Zampieri, 2021b; Vashistha and Zubiaga, 2021; Ranasinghe and Zampieri, 2021c; Nozza, 2021; Gaikwad et al., 2021; Pamungkas et al., 2021a; Aluru et al., 2021; Salaam et al., 2022; Husain and Uzuner, 2022; Gupta et al., 2022; Montariol et al., 2022; Ryzhova et al., 2022; Muti and Barrón-Cedeño, 2022; Phung and Cloos, 2022; Bigoulaeva et al., 2022; Boulouard et al., 2022; Röttger et al., 2022a; De la Peña Sarracén and Rosso, 2022; Shi et al., 2022; Deshpande et al., 2022; Eronen et al., 2022; Mozafari et al., 2022; Zia et al., 2022; Röttger et al., 2022b; Das et al., 2022; zahra El-Alami et al., 2022; Kumar and Sachdeva, 2022; Plaza-del arco et al., 2023; Zhou et al., 2023; Bigoulaeva et al., 2023; Awal et al., 2023) |
| Model Type | Model | Reference |
@@
| Machine | NB | (Ibrohim and Budi, 2019a; Burtenshaw and Kestemont, 2021; Rodríguez et al., 2021; Arango et al., 2022) |
| Learning | LR | (Aluru et al., 2020; Vashistha and Zubiaga, 2021; Aluru et al., 2021; Arango Monnar et al., 2022; Deshpande et al., 2022; Paul et al., 2023) |
| SVM | (Ibrohim and Budi, 2019a; Pamungkas and Patti, 2019; Pamungkas et al., 2020; Jiang and Zubiaga, 2021; Rodríguez et al., 2021; Bigoulaeva et al., 2021; Paul et al., 2023) | |
| other | (Hahn et al., 2021) | |
| Deep | MLP | (Jiang and Zubiaga, 2021; Pelicon et al., 2021a; Kumar and Sachdeva, 2022; Paul et al., 2023) |
@@
| Learning | CNN | (Mathur et al., 2018b; Elouali et al., 2020; Markov et al., 2021; Jiang and Zubiaga, 2021; Bigoulaeva et al., 2021, 2022; Jiang and Zubiaga, 2023; Bigoulaeva et al., 2023; Paul et al., 2023) |
@@
| LSTM | (Chiril et al., 2019; Kapoor et al., 2019; Pamungkas and Patti, 2019; Ousidhoum et al., 2019; Arango et al., 2022; Stappen et al., 2020; Pamungkas et al., 2020; Markov et al., 2021; Burtenshaw and Kestemont, 2021; Jiang and Zubiaga, 2021; Vitiugin et al., 2021; Bigoulaeva et al., 2021; Pamungkas et al., 2021a; Phung and Cloos, 2022; Bigoulaeva et al., 2022; Shi et al., 2022; Arango Monnar et al., 2022; Kumar and Sachdeva, 2022; Bigoulaeva et al., 2023; Paul et al., 2023) | |
@@
| other | (Ousidhoum et al., 2019; Bansal et al., 2020; Jiang and Zubiaga, 2021; Vashistha and Zubiaga, 2021; Arango Monnar et al., 2022; Deshpande et al., 2022; Kumar and Sachdeva, 2022; De la Peña Sarracén and Rosso, 2022) | |
@@
| Transformers | BERT variants | (Sohn and Lee, 2019; Pamungkas et al., 2020; Aluru et al., 2020; Markov et al., 2021; Arango et al., 2021; Hande et al., 2021a; Vashistha and Zubiaga, 2021; Muti and Barrón-Cedeño, 2022; Boulouard et al., 2022; Arango Monnar et al., 2022; Zia et al., 2022; zahra El-Alami et al., 2022; Plaza-del arco et al., 2023; Zhou et al., 2023; Jiang and Zubiaga, 2023; Paul et al., 2023; Husain and Uzuner, 2022; Pant and Dadu, 2020) |
@@
| mBERT | (Sohn and Lee, 2019; Glavaš et al., 2020; Aluru et al., 2020; Burtenshaw and Kestemont, 2021; Tiţa and Zubiaga, 2021; Jiang and Zubiaga, 2021; Rodríguez et al., 2021; Song et al., 2021; Pelicon et al., 2021b, a; Hande et al., 2021b; Ranasinghe and Zampieri, 2021c; Nozza, 2021; Pamungkas et al., 2021a; Aluru et al., 2021; Montariol et al., 2022; Muti and Barrón-Cedeño, 2022; Bigoulaeva et al., 2022; Boulouard et al., 2022; De la Peña Sarracén and Rosso, 2022; Deshpande et al., 2022; Eronen et al., 2022; Das et al., 2022; zahra El-Alami et al., 2022; Zhou et al., 2023; Bigoulaeva et al., 2023; Awal et al., 2023) | |
| XLM variants | (Corazza et al., 2020; Stappen et al., 2020; Hande et al., 2021a; Zhou et al., 2023; Montariol et al., 2022; Röttger et al., 2022a, b) | |
@@
| XLM-R | (Pant and Dadu, 2020; Glavaš et al., 2020; Ranasinghe and Zampieri, 2020; Firmino et al., 2021; Tiţa and Zubiaga, 2021; Jiang and Zubiaga, 2021; Song et al., 2021; Ranasinghe and Zampieri, 2021a; Hande et al., 2021a, b; Ranasinghe and Zampieri, 2021b, c; Nozza, 2021; Gaikwad et al., 2021; Salaam et al., 2022; Gupta et al., 2022; Montariol et al., 2022; Ryzhova et al., 2022; De la Peña Sarracén and Rosso, 2022; Arango Monnar et al., 2022; Eronen et al., 2022; Mozafari et al., 2022; Zia et al., 2022; Plaza-del arco et al., 2023; Awal et al., 2023) | |
@@
| other | (Rodríguez et al., 2021; Pelicon et al., 2021b; Hande et al., 2021a, b; Husain and Uzuner, 2022; Gupta et al., 2022; Montariol et al., 2022; Plaza-del arco et al., 2023; Jiang and Zubiaga, 2023; Pant and Dadu, 2020; Das et al., 2022) | |
| LLMs | FLAN-T5, mT0 | (Plaza-del arco et al., 2023) |
@@
To implement instance transfer, a correspondence between instances in the source and target languages must be created on the data level. Correspondence refers to the pair of instances with the same meaning of texts or identical labels (Pikuliak et al., 2021), e.g., parallel data with the same label matching, and translated text retains the same label of its original counterpart. Based on the established correspondence, texts or labels can be transitioned from one language to the other . Subsequently, these projected texts or labels can be used for further training and fine-tuning stages.